# Chemical Language Model Linker: blending text and molecules with modular adapters

### v1

@@ Abstract
The development of large language models and multi-modal models has enabled the appealing idea of generating novel molecules from text descriptions. Generative modeling would shift the paradigm from relying on large-scale chemical screening to find molecules with desired properties to directly generating those molecules. However, multi-modal models combining text and molecules are often trained from scratch, without leveraging existing high-quality pretrained models. That approach consumes more computational resources and prohibits model scaling. In contrast, we propose a lightweight adapter-based strategy named Chemical Language Model Linker (ChemLML). ChemLML blends the two single domain models and obtains conditional molecular generation from text descriptions while still operating in the specialized embedding spaces of the molecular domain. ChemLML can tailor diverse pretrained text models for molecule generation by training relatively few adapter parameters. We find that the choice of molecular representation used within ChemLML, SMILES versus SELFIES, has a strong influence on conditional molecular generation performance. SMILES is often preferable despite not guaranteeing valid molecules. We raise issues in using the large PubChem dataset of molecules and their associated descriptions for evaluating molecule generation and provide a filtered version of the dataset as a generation test set. To demonstrate how ChemLML could be used in practice, we generate candidate protein inhibitors and use docking to assess their quality.
@@ 1 Introduction
Machine learning methods have emerged as powerful tools in biology and chemistry (Ching et al., 2018). These methods have been applied extensively across a range of small molecule and protein problems, from molecular property prediction (Gilmer et al., 2017) and molecule interaction prediction (Deng et al., 2020) to protein structure prediction (Krishna et al., 2024) and inverse folding design (Zheng et al., 2023).
@@
Generative machine learning models are a particularly attractive area of research as an alternative to traditional experimental and computational chemical screening. Directly generating molecules with desired properties can avoid performing high-throughput chemical screens and large-scale virtual screens on molecule databases (Alnammi et al., 2023). By learning the underlying chemical patterns and structures from vast datasets, generative models such as generative adversarial networks (Kadurin et al., 2017), variational autoencoders (Jin et al., 2018), and autoregressive models (Bagal et al., 2021) have enabled the automated design of molecules with desired properties. Recent efforts in generative chemical modeling include controllable cross-modality generation. These methods include contrastive learning, multi-task training, and finetuning of large language models (LLMs).
@@
LLMs have dramatically reshaped the landscape of natural language processing (NLP), demonstrating remarkable capabilities in generating coherent and contextually appropriate text across diverse domains. Built on transformer architectures and trained on broad corpora, LLMs achieve a nuanced representation of language patterns, syntax, and semantics. This advanced representation enables them to perform a wide array of tasks, from simple text completion to complex question answering and summarizing. A critical aspect of their development has been the observed scaling laws (Kaplan et al., 2020), which suggest that the performance of these models improves predictably with an increase in model size and data volume. With LLMs’ achievements across different domains, researchers are increasingly exploring their potential to address challenges in the fields of biology and chemistry (Zhang et al., 2024).
@@
Although general purpose LLMs show some capability to operate with molecules, it is limited (Christofidellis et al., 2023; Microsoft Research AI4Science and Microsoft Azure Quantum, 2023; Mirza et al., 2024). A natural question is how to construct a multi-modal model that operates on natural language and chemicals. Previous strategies include aligning the embedding spaces between text and molecules through contrastive learning, T5-based (Raffel et al., 2020) text-to-molecule translation, training a single model on multiple molecule-related tasks, and finetuning LLMs on molecule-specific corpora. However, all these methods have varying limitations. The T5 model requires additional pretraining and cannot apply previously pretrained models on related areas. Most notably, prior work typically combines a single natural language model with a single molecular model without exploring the interactions between models and molecular representations.
@@
To address these issues, we introduce a flexible approach for merging arbitrary text-based language models with molecule generators named Chemical Language Model Linker (ChemLML). A chemical linker is a molecular structure that connects two or more molecules or functional groups via stable chemical bonds, widely used in drug development (for example, proteolysis targeting chimeras (Sakamoto et al., 2001) and fragment-based drug design (Grenier et al., 2023)), biomolecular research, and material science. We borrow the idea of “linker”, and our “linker” metaphorically connects natural and molecular languages through adapters instead of linking different molecules or functional groups. Despite the name, ChemLML is not specifically designed to generate linker molecules (Imrie et al., 2020; Guan et al., 2023; Igashov et al., 2024). ChemLML makes full use of pretrained models for text and molecules, which can greatly reduce the training time. The main contributions are:
@@
    We present ChemLML, which requires significantly fewer trainable parameters compared to models of the same scale for text-guided molecule design tasks, achieving strong performance by training adapters only.
@@
    Our case study suggests that ChemLML may be able to generate candidate inhibitors of protein targets from text descriptions alone based on its docking performance.
@@ 2 Related Work
Contrastive learning Contrastive learning has emerged as a powerful technique for learning rich, meaningful representations in multiple domains. SimCLR (Chen et al., 2020) performs contrastive learning between different augmented views of the same image to learn visual representations. MolCLR (Wang et al., 2022) extends a similar method to graph neural networks on molecule graphs. CLIP (Radford et al., 2021) introduces a novel approach to contrastive learning that unifies language and vision modalities. The model is trained to predict which images are paired with which texts. MoleculeSTM (Liu et al., 2023a) trains a multi-modal molecule-text model by learning a joint representation of molecules’ chemical structures and textual descriptions via contrastive learning. MoleculeSTM is able to perform structure-text retrieval and molecule editing tasks.
@@
Multi-task learning Multi-task learning is applied to molecular tasks because it enables models to leverage shared information across tasks, often improving generalization and performance by learning common patterns. The origins of some multi-modal text-and-chemical models came from NLP. For instance, MT-DNN (Liu et al., 2019) leverages the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) architecture as a shared encoder and simultaneously learns across multiple NLP tasks. By adding task-specific layers on top of BERT, MT-DNN integrates the advantages of shared knowledge to improve generalization and performance on those tasks. The T5 model (Raffel et al., 2020) treats all NLP tasks as text-to-text problems, enabling unified, scalable multi-task learning with a single model. The model is pretrained and finetuned on diverse tasks for enhanced performance and efficiency. MolT5 (Edwards et al., 2022) uses a T5 model to perform molecule-to-text and text-to-molecule tasks at the same time. Text+ChemT5 (Christofidellis et al., 2023) further extends the number of molecule-related tasks.
@@
Chemical language models Chemical language models are models that are trained on string representations of molecules (Grisoni, 2023). They primarily fall into two categories of architectures: BERT and generative pretrained transformer (GPT). ChemBERTa (Chithrananda et al., 2020) is a specialized adaptation of the BERT model, tailored for chemistry. It is pretrained with masked language modeling and finetuned for molecule property prediction. MolGPT is a GPT model tailored for molecular generation tasks that is trained by predicting the next token. For a more comprehensive understanding of specific models, see recent reviews (Zhang et al., 2024; Liao et al., 2024; Ramos et al., 2024).
@@
Text-guided molecule generation Text2Mol (Edwards et al., 2021) proposes the problem of generating molecules from their text descriptions, which is framed as a molecule retrieval task. Instead of simply retrieving the most similar molecules or text, MolT5 (Edwards et al., 2022) trains a T5 model that can perform text-guided molecule generation and molecule captioning tasks at the same time. Text+Chem T5 (Christofidellis et al., 2023) further trains a T5 model on multiple tasks related to molecules and improves the model’s capability on text-guided molecule generation. nach0 (Livne et al., 2024) is a multi-modal model also based on the T5 architecture that is pretrained on PubMed abstracts, patent descriptions, and ZINC chemicals. MoMu (Su et al., 2022) uses contrastive learning to bridge a graph-based molecule encoder and a natural language encoder to perform multiple molecular tasks. TGM-DLM (Gong et al., 2024) leverages diffusion models to address the limitations of autoregressive methods in molecule generation. MolXPT (Liu et al., 2023c) includes wrapped sentences in its pretraining dataset, which are sentences from scientific text that contain molecules that are detected and replaced with the matching SMILES string. MolFM (Luo et al., 2023) includes a knowledge graph over drugs, proteins, and diseases in its multi-modal model. AMAN (Zhao et al., 2024a) introduces a graph transformer for the molecule representation and trains using a combination of adversarial loss and a triplet loss that includes negative samples. ChemDFM (Zhao et al., 2024b) performs chemistry domain pretraining and instruction tuning on multiple molecule-related tasks with Llama2, developing an dialogue-based chemical LLM. Mol-Instruction (Fang et al., 2023a) prepares datasets for instruction tuning LLMs, which includes text-guided molecule generation instructions. MolReGPT (Li et al., 2024) uses in-context learning with ChatGPT and Llama 2. In-context learning with ChatGPT was also used to generate synthetic training data for molecule generation (Chen et al., 2024). MoleculeSTM (Liu et al., 2023a) can perform molecule captioning, text-based molecule generation, and text-based molecule editing. A comprehensive review presents additional approaches Pei et al. (2024). The recent Language + Molecules Workshop (Edwards et al., 2024), whose dataset includes biomedical and other chemical properties, also inspired new methods.
@@ 3.2 Preliminaries
Simplified molecular input line-entry system (SMILES) SMILES is a notation used to represent molecules as strings (Krenn et al., 2020). SMILES strings describe atomic composition, connectivity, and stereochemistry using simple rules. For example, the SMILES of 4-Methylphenol is Cc1ccc(O)cc1.
Self-referencing embedded strings (SELFIES) SELFIES is an alternative representation for molecules designed to overcome some limitations of SMILES (Krenn et al., 2020). It ensures the generation of syntactically and semantically valid molecular structures. This is achieved through a self-referencing system that encodes rules and constraints within the string itself, making SELFIES particularly useful for applications in generative models. The SELFIES of 4-Methylphenol is [C][C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][O]. It is much longer than the corresponding SMILES format.
@@ 3.3 ChemLML model architecture
Text pretrained models We used three pretrained text LLMs that are specialized for scientific text. SciBERT (Beltagy et al., 2019) is a lightweight model with 110M parameters. Galactica (Taylor et al., 2022) was reported to have the best performance in a chemistry-related evaluation (Mirza et al., 2024). Due to GPU memory constraints, we only used Galactica-125M, 1.3B, and 6.7B. We also used the T5 encoder from Text+ChemT5 (Christofidellis et al., 2023).
Molecule pretrained models We focused on pretrained models that use SMILES or SELFIES string representations. MolGPT (Bagal et al., 2021) is a GPT model. MolGen (Fang et al., 2023b) is a BART-style model that has two stages of pretraining. During the first stage, it randomly masks some of the SELFIES tokens, encodes the corrupted SELFIES using a bidirectional model, calculates the likelihood of SELFIES string with a left-to-right autoregressive decoder, and calculates the reconstruction loss. In the second stage, it introduces the domain-agnostic molecular prefix as a domain instructor to facilitate the transfer of knowledge across diverse domains. We only use the MolGen autoregressive decoder. MolXPT (Liu et al., 2023c) is another GPT-style model, which unifies text and molecules during pretraining. It replaces the molecule name with a SMILES string. We only use the molecule decoding capability of MolXPT by feeding the start-of-molecule token at the beginning during decoding.
Complete architecture We use the text pretrained model as a text encoder and the molecule pretrained model as a molecule decoder. We add a chemical adapter to the last layer of the molecule decoder that takes both text and molecule embeddings as input and outputs a molecule embedding. A similar architecture was applied in the protein domain with LM-Design (Zheng et al., 2023) and ProtT3 (Liu et al., 2024). The model architecture is shown in Fig. 1.
@@ 3.5 Metrics
We intend generated molecules to exhibit specific properties. Following the molecular similarity principle (Duran-Frigola et al., 2020), this would result in the generated molecules having structures similar to the ground truth molecules, thereby achieving a high degree of similarity and low diversity during evaluations. Therefore, we primarily focus on various similarity metrics and calculate the similarity between generated molecules and ground truth molecules. We use MACCS fingerprints, RDKit fingerprints (Landrum, 2018), and Morgan fingerprints as different molecule representations that produce different similarity scores. The similarities between these fingerprints are calculated with Tanimoto similarity. We discuss other metrics not used here in the Supplementary Methods.
SMILES-based generation will generate syntactically invalid molecules sometimes, so we also include validity as a metric. Even though SELFIES is designed to include 100% validity, the SELFIES-based model might fail to generate a valid molecule, such as samples with an [EOS] token at the beginning. The validity metric only evaluates syntactic correctness without accounting for the molecule’s 3D structure.
@@ 3.6 Datasets
  • Selected ZINC-15: Chemformer (Irwin et al., 2022) selected approximately 100M molecules from 1.5B available molecules from ZINC-15 (Sterling and Irwin, 2015) with the following constraints: reactivity set to reactive, purchasability set to annotated, molecular weight $\leq 500$ Daltons, and LogP $\leq 5$. MolGen uses this dataset during the first stage of pretraining.
  • NPASS: Natural Product Activity & Species Source Database (NPASS) (Zeng et al., 2018) is a natural product domains database, which is used in the second pretraining stage of MolGen.
  • MOSES: MOSES (Polykovskiy et al., 2020) is a cleaned version of ZINC, containing 1.9M molecules. It removes the atoms besides C, N, S, O, F, Cl, Br, H and cycles longer than 8 atoms. It does not contain chirality information. Models pretrained on the MOSES dataset, such as MolGPT, are unable to process SMILES strings that include atoms beyond this dataset’s scope or SMILES strings that contain chirality information.
  • PubChem and PubMed: MolXPT trains on the titles and abstracts from 30M PubMed entries, SMILES strings from 30M chemicals from PubChem (Kim et al., 2023), and 8M sequences in which chemical names are replaces with SMILES.
@@
Molecule description datasets We use two different molecule description datasets. We train and test all the models on ChEBI-20 and use PubChem as an additional test set. We use the training set of ChEBI-20 for training. Then, we tune the hyperparameters based on the convergence status on the ChEBI-20 training set and the performance on the ChEBI-20 validation set. Finally, we test on the ChEBI-20 and PubChem test sets.
  • ChEBI-20: The Chemical Entities of Biological Interest (ChEBI) database (Hastings et al., 2016) is a dictionary of small molecules that includes text descriptions. Edwards et al. (2021) created the ChEBI-20 dataset by combining compounds from PubChem with ChEBI annotations and retaining the 33,010 compound-description pairs for which the description had more than 20 words. The ChEBI-20 training, validation, and test set splits of 80%, 10%, and 10% are also from Edwards et al. (2021).
  • PubChem: In our evaluations, we assume that each text description is informative and associated with one unique molecule. Many examples in the complete PubChem dataset (Kim et al., 2023) do not meet that assumption, so we apply several types of filters on the molecules and text descriptions. In short, we select descriptions greater than 30 words and remove one-to-many description-molecule pairs. In this way, we remove ambiguous instances and improve the quality of our evaluations. After filtering, we remove the intersection of molecules between PubChem and ChEBI-20 dataset and obtain a dataset that has 11,563 examples. We randomly sample 3,000 examples from this dataset and use it as an additional test set for models trained on ChEBI-20. To demonstrate the impact of this filtering, we also report results on the unfiltered PubChem dataset with 2,576 molecule descriptions, which we refer to as PubChem-unfiltered. More details of PubChem dataset filtering and processing are discussed in Supplementary Methods.

### v2

@@ Abstract
The development of large language models and multi-modal models has enabled the appealing idea of generating novel molecules from text descriptions. Generative modeling would shift the paradigm from relying on large-scale chemical screening to find molecules with desired properties to directly generating those molecules. However, multi-modal models combining text and molecules are often trained from scratch, without leveraging existing high-quality pretrained models. Training from scratch consumes more computational resources and prohibits model scaling. In contrast, we propose a lightweight adapter-based strategy named Chemical Language Model Linker (ChemLML). ChemLML blends the two single domain models and obtains conditional molecular generation from text descriptions while still operating in the specialized embedding spaces of the molecular domain. ChemLML can tailor diverse pretrained text models for molecule generation by training relatively few adapter parameters. We find that the choice of molecular representation used within ChemLML, SMILES versus SELFIES, has a strong influence on conditional molecular generation performance. SMILES is often preferable despite not guaranteeing valid molecules. We raise issues in using the entire PubChem dataset of molecules and their associated descriptions for evaluating molecule generation and provide a filtered version of the dataset as a generation test set. To demonstrate how ChemLML could be used in practice, we generate candidate protein inhibitors and use docking to assess their quality and also generate candidate membrane permeable molecules.
keywords:
American Chemical Society,
[…]
@@ 1 Introduction
Machine learning methods have emerged as powerful tools in biology and chemistry 1. These methods have been applied extensively across a range of small molecule and protein problems, from molecular property prediction 2 and molecule interaction prediction 3 to protein structure prediction 4 and inverse folding design 5.
@@
Generative machine learning models are a particularly attractive area of research as an alternative to traditional experimental and computational chemical screening. Directly generating molecules with desired properties can avoid performing high-throughput chemical screens and large-scale virtual screens on molecule databases 6. By learning the underlying chemical patterns and structures from vast datasets, generative models such as generative adversarial networks 7, variational autoencoders 8, and autoregressive models 9 have enabled the automated design of molecules with desired properties. Recent efforts in generative chemical modeling include controllable cross-modality generation. These methods include contrastive learning, multi-task training, and finetuning of large language models (LLMs).
@@
LLMs have dramatically reshaped the landscape of natural language processing (NLP), demonstrating remarkable capabilities in generating coherent and contextually appropriate text across diverse domains. Built on transformer architectures and trained on broad corpora, LLMs achieve a nuanced representation of language patterns, syntax, and semantics. This advanced representation enables them to perform a wide array of tasks, from simple text completion to complex question answering and summarizing. A critical aspect of their development has been the observed scaling laws 10, which suggest that the performance of these models improves predictably with an increase in model size and data volume. With LLMs’ achievements across different domains, researchers are increasingly exploring their potential to address challenges in the fields of biology and chemistry 11.
@@
Although general purpose LLMs show some capability to operate with molecules, it is limited 12, 13, 14. A natural question is how to construct a multi-modal model that operates on natural language and chemicals. Previous strategies include aligning the embedding spaces between text and molecules through contrastive learning, T5-based 15 text-to-molecule translation, training a single model on multiple molecule-related tasks, and finetuning LLMs on molecule-specific corpora. However, all these methods have varying limitations. The T5 model requires additional pretraining and cannot apply previously pretrained models on related areas. Most notably, prior work typically combines a single natural language model with a single molecular model without exploring the interactions between models and molecular representations.
@@
To address these issues, we introduce a flexible approach for merging arbitrary text-based language models with molecule generators named Chemical Language Model Linker (ChemLML). A chemical linker is a molecular structure that connects two or more molecules or functional groups via stable chemical bonds, widely used in drug development (for example, proteolysis targeting chimeras 16 and fragment-based drug design 17), biomolecular research, and material science. We borrow the idea of “linker”, and our “linker” metaphorically connects natural and molecular languages through adapters instead of linking different molecules or functional groups. Despite the name, ChemLML is not specifically designed to generate linker molecules 18, 19, 20. ChemLML makes full use of pretrained models for text and molecules, which can greatly reduce the training time. The main contributions are:
@@
    We present ChemLML, which requires far fewer trainable parameters compared to models of the same scale for text-guided molecule design tasks, achieving strong performance by training adapters only.
@@
    Our case studies suggest that ChemLML may be able to generate candidate inhibitors of protein targets based on its docking performance and membrane permeable molecules from text descriptions alone.
@@ 2 Related Work
Contrastive learning Contrastive learning has emerged as a powerful technique for learning rich, meaningful representations in multiple domains. SimCLR 21 performs contrastive learning between different augmented views of the same image to learn visual representations. MolCLR 22 extends a similar method to graph neural networks on molecule graphs. CLIP 23 introduces an approach to contrastive learning that unifies language and vision modalities. The model is trained to predict which images are paired with which texts. MoleculeSTM 24 trains a multi-modal molecule-text model by learning a joint representation of molecules’ chemical structures and textual descriptions via contrastive learning. MoleculeSTM is able to perform structure-text retrieval and molecule editing tasks. CLAMP 25 utilizes contrastive learning to associate molecules and their corresponding assay descriptions, thus enhancing the model’s capability in few-shot and zero-shot molecule activity prediction tasks.
@@
Multi-task learning Multi-task learning is applied to molecular prediction tasks because it enables models to leverage shared information across tasks, often improving generalization and performance by learning common patterns. The origins of some multi-modal text-and-chemical models came from NLP. For instance, MT-DNN 26 leverages the Bidirectional Encoder Representations from Transformers (BERT) 27 architecture as a shared encoder and simultaneously learns across multiple NLP tasks. By adding task-specific layers on top of BERT, MT-DNN integrates the advantages of shared knowledge to improve generalization and performance on those tasks. The T5 model 15 treats all NLP tasks as text-to-text problems, enabling unified, scalable multi-task learning with a single model. The model is pretrained and finetuned on diverse tasks for enhanced performance and efficiency. MolT5 28 uses a T5 model to perform molecule-to-text and text-to-molecule tasks at the same time. Text+ChemT5 12 further extends the number of molecule-related tasks.
@@
Chemical language models Chemical language models are models that are trained on string representations of molecules 29. They primarily fall into two categories of architectures: BERT and generative pretrained transformer (GPT). ChemBERTa 30 is a specialized adaptation of the BERT model, tailored for chemistry. It is pretrained with masked language modeling and finetuned for molecule property prediction. MolGPT is a GPT model tailored for molecular generation tasks that is trained by predicting the next token. For a more comprehensive understanding of specific models, see recent reviews 11, 31, 32.
@@
Text-guided molecule generation Text2Mol 33 proposes the problem of generating molecules from their text descriptions, which is framed as a molecule retrieval task. MolT5 28 and Text+Chem T5 12 mentioned previously are examples of models that can generate molecules given text inputs. nach0 34 is a multi-modal model also based on the T5 architecture that is pretrained on PubMed abstracts, patent descriptions, and ZINC chemicals. MoMu 35 uses contrastive learning to bridge a graph-based molecule encoder and a natural language encoder to perform multiple molecular tasks. TGM-DLM 36 leverages diffusion models to address the limitations of autoregressive methods in molecule generation. MolXPT 37 includes wrapped sentences in its pretraining dataset, which are sentences from scientific text that contain molecules that are detected and replaced with the matching SMILES string. MolFM 38 includes a knowledge graph over drugs, proteins, and diseases in its multi-modal model. AMAN 39 introduces a graph transformer for the molecule representation and trains using a combination of adversarial loss and a triplet loss that includes negative samples. ChemDFM 40 performs chemistry domain pretraining and instruction tuning on multiple molecule-related tasks with Llama2, developing an dialogue-based chemical LLM. Mol-Instruction 41 prepares datasets for instruction tuning LLMs, which includes text-guided molecule generation instructions. MolReGPT 42 uses in-context learning with ChatGPT and Llama 2. In-context learning with ChatGPT can also generate synthetic training data for molecule generation 43. MoleculeSTM 24 can perform molecule captioning, text-based molecule generation, and text-based molecule editing. The recent Language + Molecules Workshop 44, whose dataset includes biomedical and other chemical properties, also inspired new methods. A comprehensive review presents additional approaches 45.
@@ 3.2 Preliminaries
Simplified molecular input line-entry system (SMILES) SMILES is a notation used to represent molecules as strings 46. SMILES strings describe atomic composition, connectivity, and stereochemistry using simple rules. For example, the SMILES of 4-methylphenol is Cc1ccc(O)cc1.
Self-referencing embedded strings (SELFIES) SELFIES is an alternative representation for molecules designed to overcome some limitations of SMILES 46. It ensures the generation of syntactically and semantically valid molecular structures. This is achieved through a self-referencing system that encodes rules and constraints within the string itself, making SELFIES particularly useful for applications in generative models. The SELFIES of 4-methylphenol is [C][C][=C][C][=C][Branch1][Branch1][C][=C][Ring1][=Branch1][O]. It is much longer than the corresponding SMILES format.
@@ 3.3 ChemLML model architecture
Text pretrained models We used three pretrained text LLMs that are specialized for scientific text. SciBERT 47 is a lightweight model with 110M parameters. Galactica 48 was reported to have the best performance in a chemistry-related evaluation 14. Although Chartier-Edwards et al. 49 pointed out hallucination and performance issues with Galactica, we do not use Galactica’s generative capabilities. We use it as an embedding model and explore whether LLMs offer improved scientific text representation capabilities over time in comparison to older models like SciBERT. Due to GPU memory constraints, we only used Galactica-125M, 1.3B, and 6.7B. Finally, we also used the T5 encoder from Text+ChemT5 12.
Molecule pretrained models We focused on pretrained models that use SMILES or SELFIES string representations. MolGPT 9 is a GPT model. MolGen 50 is a BART-style model that has two stages of pretraining. During the first stage, it randomly masks some of the SELFIES tokens, encodes the corrupted SELFIES using a bidirectional model, calculates the likelihood of the SELFIES string with a left-to-right autoregressive decoder, and calculates the reconstruction loss. In the second stage, it introduces the domain-agnostic molecular prefix as a domain instructor to facilitate the transfer of knowledge across diverse domains. We only use the MolGen autoregressive decoder. MolXPT 37 is another GPT-style model, which unifies text and molecules during pretraining. It replaces the molecule name with a SMILES string. We only use the molecule decoding capability of MolXPT by feeding the start-of-molecule token at the beginning during decoding.
Complete architecture We use the pretrained text model as a text encoder and the pretrained molecule model as a molecule decoder. We add a chemical adapter to the last layer of the molecule decoder that takes both text and molecule embeddings as input and outputs a molecule embedding. A similar architecture was applied in the protein domain with LM-Design 5 and ProtT3 51. The model architecture is shown in Figure 1.
@@ 3.5 Metrics
We intend generated molecules to exhibit specific properties. Following the molecular similarity principle 52, this would result in the generated molecules having structures similar to the ground truth molecules, thereby achieving a high degree of similarity and low diversity during evaluations. Therefore, we primarily focus on various similarity metrics and calculate the similarity between generated molecules and ground truth molecules. We use MACCS fingerprints, RDKit fingerprints 53, and Morgan fingerprints as different molecule representations that produce different similarity scores. The similarities between these fingerprints are calculated with Tanimoto similarity. We discuss other metrics not used here in the Supplementary Methods.
SMILES-based generation sometimes generates syntactically invalid molecules, so we also include validity as a metric. Although valid SELFIES strings are guaranteed to be syntactically valid chemicals, the SELFIES-based model does not always generate valid SELFIES strings, such as output strings with an [EOS] token at the beginning. The validity metric only evaluates syntactic correctness without accounting for the molecule’s 3D structure.
@@ 3.6 Datasets
  • Selected ZINC-15: Chemformer 54 selected approximately 100M molecules from 1.5B available molecules from ZINC-15 55 with the following constraints: reactivity set to reactive, purchasability set to annotated, molecular weight $\leq 500$ Daltons, and LogP $\leq 5$. MolGen uses this dataset during the first stage of pretraining.
  • NPASS: Natural Product Activity & Species Source Database (NPASS) 56 is a natural product domains database, which is used in the second pretraining stage of MolGen.
  • MOSES: MOSES 57 is a cleaned version of ZINC, containing 1.9M molecules. It removes the atoms besides C, N, S, O, F, Cl, Br, and H and cycles longer than 8 atoms. It does not contain chirality information. Models pretrained on the MOSES dataset, such as MolGPT, are unable to process SMILES strings that include atoms beyond this dataset’s scope or SMILES strings that contain chirality information.
  • PubChem and PubMed: MolXPT trains on the titles and abstracts from 30M PubMed entries, SMILES strings from 30M chemicals from PubChem 58, and 8M sequences in which chemical names are replaces with SMILES.
@@
Molecule description datasets We use two different molecule description datasets. We train and test all the models on ChEBI-20 and use PubChem as an additional test set. We use the ChEBI-20 training set for training. Then, we tune the hyperparameters based on the convergence status on the ChEBI-20 training set and the performance on the ChEBI-20 validation set. Finally, we test on the ChEBI-20 and PubChem test sets.
  • ChEBI-20: The Chemical Entities of Biological Interest (ChEBI) database 59 is a dictionary of small molecules paired with text descriptions. Edwards et al. 33 created the ChEBI-20 dataset by combining PubChem compounds with ChEBI annotations and retaining the 33,010 compound-description pairs for which the description had more than 20 words. The ChEBI-20 training, validation, and test set splits of 80%, 10%, and 10% are also from Edwards et al. 33.
  • PubChem: In our evaluations, we assume that each text description is informative and associated with one unique molecule. Many examples in the complete PubChem dataset 58 do not meet that assumption, so we apply several types of filters on the molecules and text descriptions. In short, we select descriptions greater than 30 words and remove one-to-many description-molecule pairs. In this way, we remove ambiguous instances and improve the quality of our evaluations. After filtering, we remove the intersection of molecules between PubChem and ChEBI-20 dataset and obtain a dataset that has 11,563 examples. We randomly sample 3,000 examples from this dataset, which we refer to as PubChem-filtered, and use it as an additional test set for models trained on ChEBI-20. To demonstrate the impact of this filtering, we also report results on 2,576 examples sampled from the unfiltered PubChem dataset, which we refer to as PubChem-unfiltered. More details of PubChem dataset filtering and processing are discussed in Supplementary Methods.